OpenStudy (allgemeine-klause01):

The slope-intercept form of the equation of a line that passes through point (–3, 8) is y = –x + 6. What is the point-slope form of the equation for this line?

pooja195 (pooja195):

y=mx+b <----slope intercept form y=-x+6 <----what is the slope in the equation? m=slope

OpenStudy (allgemeine-klause01):

srry the code sucks its supposed to be y= -2/3x+6

pooja195 (pooja195):

y=mx+b <----slope intercept form y=-2/3x+6 <----what is the slope in the equation? m=slope

OpenStudy (allgemeine-klause01):

-2

pooja195 (pooja195):

keep in mind that -2/3 is a fraction it belongs together

OpenStudy (allgemeine-klause01):

oh -2/3

pooja195 (pooja195):

This is point slope form \[\huge y-y_1=m(x-x_1)\] where m=-2/3 <---we know this from the slope intercept problem (x1,y1) (-3,8) x1=? y1=?

OpenStudy (allgemeine-klause01):

x1=-3 y1=8

pooja195 (pooja195):

Good now plug those numbers where they belong into the formula! :) \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

We know that: x1=-3 y1=8 m=-2/3

OpenStudy (allgemeine-klause01):

umm y-8=-2/3(x+3)

OpenStudy (allgemeine-klause01):

?

pooja195 (pooja195):

Perfect :D

OpenStudy (allgemeine-klause01):

thx i get confused when it comes to math

pooja195 (pooja195):

Its alright practice makes perfect B)